OpenStudy (anonymous):

how do you write 2/3x = -5/2y + 2 in standard form?

OpenStudy (anonymous):

\[\frac{2}{3}x=-\frac{5}{2}y+2\] Standard form is, \[Ax+By+C=0\]\[\frac{2}{3}x+\frac{5}{2}y+(-2)=0\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No worries.

OpenStudy (anonymous):

can you convert it?

OpenStudy (anonymous):

Convert it to what?